how many grams of CoCl2 are required to produce 5.30 x 10^23 molecules of chlorine
1
Expert's answer
2020-12-08T05:32:35-0500

"n = N\/N_A = 5.30*10^{23}\/6.02*10^{23} = 0.88 mol"

"n(CoCl_2) = n(Cl_2)"

"m(CoCl_2) = nM = 0.88*130 = 114.4 g"